Can someone show and explain to me how to solve for graphs using Algebra? I need to complete the following:
These two functions: f(x) = -2x + 5 and g(x) = x^2 + 2x + 6 intersect at two points. Determine using Algebra the x and y co-ordinates of the two points correct to 5 decimal places.
Help with explanations and steps would be great.
Thanks.

Answer by scott8148(6628)

the two functions are equal at the intersection points

-2x + 5 = x^2 + 2x + 6

0 = x^2 + 4x + 1

use the quadratic formula to find the two values of x

then plug the x-values into a function (I'd use the simpler one) to get the corresponding y-values